TritanCarsons:

10x2-5+245(3x33-100)45-12

TritanCarsons:

10x2-5+245(3x33-100)45-12+1/2 /12

ILOVESPAGHETTI:

Alright, so for this problem Pemdas is going to help a lot. Some things to note, if a number in on the parenthesis then it means multiply, except when I use it to differante So first, Parenthesis So if you do the things in the parenthesis 3x33 is 99. So 99-100= -1(negative 1) Now that we figured that out, 245(-1) is -245. Our problem now looks like 10x2-5+(-245)45-12+1/2 /12 10 times 2 is 20, and that subtracted by 5 is 15. 15+(-245)= -230 -230(not specified operation?)45-12 so do 45-12=33 and whatever operation you forgot to put in the equation. also, for 1/2 and /12 I can't provide a answer, because no operation is stated. Please let me know if I messed up/did anything wrong.

ILOVESPAGHETTI:

differentiate*

ILOVESPAGHETTI:

@ilovespaghetti wrote:
Alright, so for this problem Pemdas is going to help a lot. Some things to note, if a number in on the parenthesis then it means multiply, except when I use it to differante So first, Parenthesis So if you do the things in the parenthesis 3x33 is 99. So 99-100= -1(negative 1) Now that we figured that out, 245(-1) is -245. Our problem now looks like 10x2-5+(-245)45-12+1/2 /12 10 times 2 is 20, and that subtracted by 5 is 15. 15+(-245)= -230 -230(not specified operation?)45-12 so do 45-12=33 and whatever operation you forgot to put in the equation. also, for 1/2 and /12 I can't provide a answer, because no operation is stated. Please let me know if I messed up/did anything wrong.
Also continuing with the problem, turn your fractions into decimals by dividing the numerator by the denominator. In this case, 1/2 would be 0.50. so 33+0.50 is 33.50. I can't help further with this problem

urlowkey:

Thank you so much!!!!

ILOVESPAGHETTI:

@urlowkey wrote:
Thank you so much!!!!
No problem.

urlowkey:

Why u keep sayin Blocked Message?

urlowkey:

@urlowkey wrote:
Why u keep sayin Blocked Message?